A boat travels 500 miles downstream and back. The trip downstream took 25 hours. The trip back upstream took 50 hours. What is the speed of the boat? What is the speed of the current in the river?
Answer by mananth(16949)

Boat speed =x mph
current speed =y mph
against current x-y 50.00 hours
with current x+y 20.00 hours

Distance = same= 500 km
t=d/r
500 / ( x - y )= 50.00
50 ( x - y ) = 500.00
50 x -50 y = 500 ....................1
500 / ( x + y )= 20.00
20.00 ( x + y ) = 500
20.00 x + 20.00 y = 500 ...............2
Multiply (1) by 2
Multiply (2) by 5
we get
100 x -100 y = 1000
100 x + 100 y = 2500
200 x = 3500
/ 200
x = 17.5 mph

plug value of x in (1)
50 x -50 y = 500
875 -50 y = 500
-50 y = 500 -875
-50 y = -375
y = 7.5
The speed of Boat = 17.5 mph
The speed of current = 7.5 mph
m.ananth@hotmail.ca